OpenStudy (angelwings996):

Simplify each sum or difference. State any restrictions on the variables.

OpenStudy (angelwings996):

\[\frac{ 5y+2 }{ xy ^{2} } + \frac{ 2x-4 }{ 4xy }\]

OpenStudy (phi):

the restriction is you are not allowed to divide by zero.

OpenStudy (phi):

I would look for a common denominator

hero (hero):

Multiply the first fraction by 4/4 Multiply the second fraction by y/y Afterwards, add the fractions together.

OpenStudy (angelwings996):

Okay thank you!

OpenStudy (angelwings996):

@Hero when I multiply them together, will they look like this?\[\frac{ 20y+8 }{ 4xy ^{2} } + \frac{ 2xy-4y }{ 4xy ^{2} }\]

hero (hero):

Yes, great job.

OpenStudy (angelwings996):

@Hero Alright, well I am not sure if this is right but when I added them together I got \[\frac{ 2xy+16y+8 }{ 4xy ^{2} }\]

hero (hero):

That's correct

OpenStudy (angelwings996):

Oh okay, what do I do now?

OpenStudy (phi):

State any restrictions on the variables.

OpenStudy (angelwings996):

But in the back of the book it says that the answer is something completly different

OpenStudy (phi):

you can factor out a 2 from the top and bottom. they will cancel

hero (hero):

^What he said

hero (hero):

Always look to see if the fraction can be simplified after adding

hero (hero):

Hmm, strange. I'm not sure how they got that.